OpenStudy (inkyvoyd):

Waves: does this formula look correct for the superposition of two waves of the same amplitude? How do I derive it? My two functions are sinusoids with equations \(sin(k_1*x-*\omega_1*t)\) and \(sin(k_2*x-*\omega_2*t)\)

OpenStudy (inkyvoyd):

http://prntscr.com/6pzsvv or if you prefer \(2\sin (\frac{k_2+k_1}{2}x-\frac{\omega_2+\omega_1}{2}t)\cos (\frac{k_2-k_1}{2}x-\frac{\omega_2-\omega_1}{2}t)\)

OpenStudy (inkyvoyd):

oh think I found it http://web.clark.edu/ggrey/Physics102/Notes/superposition_of_waves.htm

OpenStudy (inkyvoyd):

forgot to say that my v is the same in magnitude for both waves

OpenStudy (michele_laino):

we can start from here: \[\begin{gathered} {\psi _1} = A\exp \left[ {i\left( {{k_1}x - {\omega _1}t} \right)} \right] \hfill \\ {\psi _2} = A\exp \left[ {i\left( {{k_2}x - {\omega _2}t} \right)} \right] \hfill \\ \end{gathered} \] and then we have to compute this: \[\left( {{\psi _1} + {\psi _2}} \right)\left( {{\psi _1} + {\psi _2}} \right)*\] where \[\psi *\] is the complex conjugate of \[\psi \]

OpenStudy (irishboy123):

@Michele_Laino +1 i was thinking that too.

OpenStudy (irishboy123):

but looks quite messy, ie isn't really a shortcut.....?

OpenStudy (inkyvoyd):

@Michele_Laino what is the meaning of having the star conjugate symbol outside of the parentheses?

OpenStudy (michele_laino):

I mean this: \[\left( {{\psi _1} + {\psi _2}} \right)* = {\psi _1}* + {\psi _2}*\]

OpenStudy (inkyvoyd):

okay thank you.

OpenStudy (michele_laino):

thank you!